OpenStudy (mos1635):

thermodynamics assume ideal gas how to calculate deferance of enthalpy betwen two states in a isothermic change?

OpenStudy (mos1635):

@experimentX that mast not be so enthalpy=internal energy +P*V H=U+P*V

OpenStudy (anonymous):

\[\Delta H =\Delta U +\Delta PV\] It's isothermic so \[\Delta U=0\]

OpenStudy (experimentx):

during isothermal change, the internal energy of gas remains constant ... well that is my assumption. Haven't been doing these stuff for long ... so not really sure.

OpenStudy (mos1635):

that is exactly my problem isothermic => P*V=constant there for ΔPV=0

OpenStudy (experimentx):

I think isothermic means same temperature.

OpenStudy (anonymous):

\[\Delta nrt=nR \Delta T =0\]

OpenStudy (mos1635):

isothermic =isothermal

OpenStudy (mos1635):

so it is ΔH=0 ?

OpenStudy (experimentx):

I think internal energy of gas depends on temperature.

OpenStudy (experimentx):

if we are keeping the temperature same as well as applying heat or absorbing heat, then of course the gas must be allowed to expand. Work is done on the gas or done by the gas.

OpenStudy (mos1635):

if so that rases next question  is ΔHa1b=ΔH12b=0 ??